# Perfect state transfer between real pure states

### v3

@@ 12 Sensitivity with respect to readout time
$\frac{d^{k}f}{dt^{k}}\bigg{|}_{\tau}=\begin{cases}(-1)^{\frac{k\mod 4}{2}}\displaystyle\sum_{j=0}^{k}(-1)^{j}{{k}\choose{j}}(\mathbf{y}^{\top}M^{j}\mathbf{y})(\mathbf{y}^{\top}M^{k-j}\mathbf{y})&{\rm{if}}\ k\ \rm{is\ even}\\ 0&{\rm{if}}\ k\ \rm{is\ odd}.\end{cases}$ (13)
@@
There is practical interest in determining how large or small $\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}$ can be, since that will give intuition on what features govern the sensitivity of the fidelity with respect to the readout time. Taking $k=2$ in (13) yields
$\begin{split}\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}=-2(\mathbf{y}^{\top}M^{2}\mathbf{y}-(\mathbf{y}^{\top}M\mathbf{y})^{2}).\end{split}$ (14)
Let $\sigma_{\mathbf{y}}(M)=\{\lambda_{1},\ldots,\lambda_{\ell}\}$ and write $\mathbf{y}=\sum_{j=1}^{\ell}c_{j}E_{\lambda_{j}}\mathbf{y}$ for some $c_{1},\ldots,c_{\ell}\in\mathbb{R}$ with $\sum_{j=1}^{\ell}c_{j}^{2}=1$. Then
$\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}=2((\mathbf{y}^{\top}M\mathbf{y})^{2}-\mathbf{y}^{\top}M^{2}\mathbf{y})=2((\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}),$
where $a_{j}=c_{j}^{2}$ for each $j=1,\ldots,\ell.$ Observe that from the Cauchy-Schwarz inequality,
$(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}=(\sum_{j=1}^{\ell}(c_{j}\lambda_{j})c_{j})^{2}\leq(\sum_{j=1}^{\ell}c_{j}^{2}\lambda_{j}^{2})(\sum_{j=1}^{\ell}c_{j}^{2})=\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
@@
$\begin{split}(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}&=(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+\epsilon(\lambda_{j_{1}}-\lambda_{j_{2}})(2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{2}}))+\epsilon^{2}(\lambda_{j_{1}}-\lambda_{j_{2}})^{2}.\end{split}$ (15)
@@
$(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}<(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
@@
we find that the minimum value for $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$ is given by $-\frac{1}{2}(\lambda_{\max}-\lambda_{\min})^{2}$, which is attained by $\mathbf{x}=\frac{1}{\sqrt{2}}(\mathbf{u}+\mathbf{v})$ and $\mathbf{y}=\frac{1}{\sqrt{2}}(\mathbf{u}-\mathbf{v}),$ where $\mathbf{u},\mathbf{v}$ are unit eigenvectors corresponding to $\lambda_{\max}$ and $\lambda_{\min},$ respectively. We summarize the above discussion as follows.
@@
Consider the Petersen graph with adjacency matrix $A,$ which has eigenvalues $3,1,-2$. Let $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$ be unit eigenvectors corresponding to $3,1$ and $-2,$ respectively and form $\mathbf{x}\in\mathbb{R}^{10}$ as $\mathbf{x}=c_{1}\mathbf{v}_{1}+c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$ where $c_{1},c_{2},c_{3}\in\mathbb{R}\backslash\{0\},c_{1}^{2}+c_{2}^{2}+c_{3}^{2}=1.$ Setting $\mathbf{y}=-c_{1}\mathbf{v}_{1}-c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$, we find that there is PST from $\mathbf{x}\mathbf{x}^{\top}$ to $\mathbf{y}\mathbf{y}^{\top}$ at time $\pi.$ According to Theorem 61, we have $0>\frac{d^{2}f}{dt^{2}}\big{|}_{\pi}\geq-\frac{25}{2}$.
Combining Theorem 61 with the proof of Corollary 59, we obtain the connected unweighted $n$-vertex graphs and the unit vectors admitting PST that attain the least value of $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$. These unit vectors are precisely those that minimize the PST time amongst all connected unweighted $n$-vertex graphs.
@@
Amongst all connected unweighted $n$-vertex graphs that admit perfect state transfer between unit vectors at time $\tau$ relative to $L$ (respectively, relative to $A$), the least value of $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$ is attained by the join graph in Theorem 58(1) (respectively, Theorem 58(2)) and unit vectors $\mathbf{x}$ and $\mathbf{y}$, where $\mathbf{y}$ is given in Corollary 59(1) (respectively, Corollary 59(2)).

### v4

@@ 12 Sensitivity with respect to readout time
$\frac{d^{k}f}{dt^{k}}\Big{|}_{\tau}=\begin{cases}(-1)^{\frac{k\mod 4}{2}}\displaystyle\sum_{j=0}^{k}(-1)^{j}{{k}\choose{j}}(\mathbf{y}^{\top}M^{j}\mathbf{y})(\mathbf{y}^{\top}M^{k-j}\mathbf{y})&{\rm{if}}\ k\ \rm{is\ even}\\ 0&{\rm{if}}\ k\ \rm{is\ odd}.\end{cases}$ (13)
@@
There is practical interest in determining how large or small $\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}$ can be, since that will give intuition on what features govern the sensitivity of the fidelity with respect to the readout time. Taking $k=2$ in (13) yields
$\begin{split}\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}=-2(\mathbf{y}^{\top}M^{2}\mathbf{y}-(\mathbf{y}^{\top}M\mathbf{y})^{2}).\end{split}$ (14)
Let $\sigma_{\mathbf{y}}(M)=\{\lambda_{1},\ldots,\lambda_{\ell}\}$ and write $\mathbf{y}=\sum_{j=1}^{\ell}c_{j}\mathbf{v}_{j}$ where each $\mathbf{v}_{j}$ is a unit vector associated with $\lambda_{j}$, each scalar $\0c_{j}\in\mathbb{R}\backslash\{0\}$ and $\sum_{j=1}^{\ell}c_{j}^{2}=1$. Then
$\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}=2((\mathbf{y}^{\top}M\mathbf{y})^{2}-\mathbf{y}^{\top}M^{2}\mathbf{y})=2\left(\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}\right),$
where $a_{j}=c_{j}^{2}$ for each $j=1,\ldots,\ell.$ Observe that from the Cauchy-Scwarz inequality,
$\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)^{2}=\left(\sum_{j=1}^{\ell}(c_{j}\lambda_{j})c_{j}\right)^{2}\leq\left(\sum_{j=1}^{\ell}c_{j}^{2}\lambda_{j}^{2}\right)\left(\sum_{j=1}^{\ell}c_{j}^{2}\right)=\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
@@
$\begin{split}\left(\sum_{j=1}^{\ell}b_{j}\lambda_{j}\right)^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}&=\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+\epsilon(\lambda_{j_{1}}-\lambda_{j_{2}})\left(2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{2}})\right)+\epsilon^{2}(\lambda_{j_{1}}-\lambda_{j_{2}})^{2}.\end{split}$ (15)
@@
$\left(\sum_{j=1}^{\ell}b_{j}\lambda_{j}\right)^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}<\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
@@
we find that the minimum value for $\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}$ is given by $-\frac{1}{2}(\lambda_{\max}-\lambda_{\min})^{2}$, which is attained by $\mathbf{x}=\frac{1}{\sqrt{2}}(\mathbf{u}+\mathbf{v})$ and $\mathbf{y}=\frac{1}{\sqrt{2}}(\mathbf{u}-\mathbf{v}),$ where $\mathbf{u},\mathbf{v}$ are unit eigenvectors corresponding to $\lambda_{\max}$ and $\lambda_{\min},$ respectively.
We summarize the above discussion as follows.
@@
Consider the Petersen graph with adjacency matrix $A,$ which has eigenvalues $3,1,-2$. Let $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$ be unit eigenvectors corresponding to $3,1$ and $-2,$ respectively and form $\mathbf{x}\in\mathbb{R}^{10}$ as $\mathbf{x}=c_{1}\mathbf{v}_{1}+c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$ where $c_{1},c_{2},c_{3}\in\mathbb{R}\backslash\{0\},c_{1}^{2}+c_{2}^{2}+c_{3}^{2}=1.$ Setting $\mathbf{y}=-c_{1}\mathbf{v}_{1}-c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$, we find that there is PST from $\mathbf{x}\mathbf{x}^{\top}$ to $\mathbf{y}\mathbf{y}^{\top}$ at time $\pi.$ According to Theorem 61, $\frac{d^{2}f}{dt^{2}}\Big{|}_{\pi}$ is bounded above by 00 and below by $-\frac{25}{2}.$
Combining Theorem 61 with the proof of Corollary 59, we obtain the connected unweighted $n$-vertex graphs and the unit vectors admitting PST that attain the least value of $\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}$. These unit vectors have the most sensitive fidelity with respect to the PST time, and these are precisely those that minimize the PST time amongst all connected unweighted $n$-vertex graphs.
@@
Amongst all connected unweighted $n$-vertex graphs that admit perfect state transfer between unit vectors at time $\tau$ relative to $L$ (respectively, relative to $A$), the least value of $\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}$ is attained by the join graph in Theorem 58(1) (respectively, Theorem 58(2)) and unit vectors $\mathbf{x}$ and $\mathbf{y}$, where $\mathbf{y}$ is given in Corollary 59(1) (respectively, Corollary 59(2)).